OpenStudy (anonymous):

2.Write the equation of the line which passes through (2, 1) and is perpendicular to x = –2. (1 point)

OpenStudy (anonymous):

Help

OpenStudy (anonymous):

the equation of a line passing through a pont(x1,y1) and having slope m is (y-y1)=m(x-x1)

OpenStudy (da_scienceman):

m really trying to knw if the answer is x = -2! No info cos both parallel lines got same gradient. But then a fool might write down y - 5 = m(x + 2) , but then m is the gradient but should be same for both parallel lines. Again, is d gradient of a parallel line nt 0?

OpenStudy (anonymous):

Yeah, I got that part but I keep getting, y = 2x -3

OpenStudy (da_scienceman):

how did u get d gradient

OpenStudy (anonymous):

Is the supposed to be 1/2? as the slope

OpenStudy (da_scienceman):

why 1/2? m nt sure if its 1/2 though! its long i dealt with this type of stuff lol

OpenStudy (anonymous):

the line given is a vertical line. so it's slope is infinity.. the line perpendicular to it must have m=0

OpenStudy (anonymous):

so here m=0

OpenStudy (da_scienceman):

recall that gradient is m = (y1 - y2)/(x1 - x2 )

OpenStudy (anonymous):

I think it's 1/2, I'm 90% certain.

OpenStudy (anonymous):

y=1?

OpenStudy (da_scienceman):

Nope on the contrary, the product of the gradietn of two perpendicular lines is -1!

OpenStudy (da_scienceman):

so if d line was L1, andother is L2, then their gradients are related by L1.L2 = -1

OpenStudy (anonymous):

Wth I'm confused you guys.

OpenStudy (anonymous):

@Katiebae wat's the ans....have u any options?

OpenStudy (da_scienceman):

i ll work it out before i get home, challenging!

OpenStudy (anonymous):

I'm sorrry but you guys arent helping

OpenStudy (da_scienceman):

r u sure ur question is correct

OpenStudy (anonymous):

I think it's y = 1/2x

OpenStudy (anonymous):

Yes it's correct^

OpenStudy (da_scienceman):

why?

OpenStudy (anonymous):

Because the y intercept becomes a zero

OpenStudy (anonymous):